Circular No. 6660

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 S. G. Djorgovski, M. R. Metzger, S. R. Kulkarni, S. C. Odewahn,

R. R. Gal, and M. A. Pahre, California Institute of Technology; D. A. Frail, National Radio Astronomy Observatory; E. Costa and M. Feroci, and the rest of the BeppoSAX team write: “Observations at Palomar indicate that the optical counterpart of GRB 970508 is continuing to decline in brightness; on May 13.18 UT, its Gunn r-band magnitude was 20.76 ± 0.15. We also detect a faint blue galaxy with r = 24.8 and g = 24.4, located ∼ 4′′.3 east and 3′′.5 north of the variable. The two may be associated, and/or the galaxy may be responsible for the absorption lines seen in the spectrum of the object at z = 0.835 (IAUC 6655).”

C. Kouveliotou, Universities Space Research Association (USRA); M. S. Briggs and R. Preece, University of Alabama in Huntsville (UAH); and G. J. Fishman, C. A. Meegan, and B. A. Harmon, Marshall Space Flight Center, NASA, report on behalf of the CGRO BATSE team: “GRB 970508 is a weak, classical γ-ray burst in our data, consisting of a single pulse with duration ∼ 3.6 s (FWHM) and total duration ∼ 35 s. The spectrum of the event is consistent with a single power law with photon index −1.5 ± 0.05 between 20 and 1000 keV. Fits with more complex models indicate a 2σ lower limit for the energy, where the burst emits most of its power (E_{peak}) at ∼ 500 keV. The estimated total fluence (20–1000 keV) is (3.1 ± 0.2) × 10^{−6} erg cm^{−2}, and its peak flux (50–300 keV) integrated over 1.024 s is (1.66 ± 0.06) × 10^{−7} erg cm^{−2} s. This peak flux corresponds to an isotropic luminosity of ∼ 6 × 10^{50} erg/s for the reported redshift of 0.835 (IAUC 6655), which agrees with fits to the γ-ray-burst intensity distribution. An earth-occultation search on May 8 and 9 at the location of the x-ray source (IAUC 6656) reveals no emission up to a 1-day 3σ limit of ∼ 30 mCrab (20–100 keV), assuming the above power-law spectral shape.”

P. J. Groot and T. J. Galama, University of Amsterdam (UoA); J. van Paradijs, UoA and UAH; C. Kouveliotou, USRA; M. Centurion and J. Telting, Isaac Newton Group of Telescopes; P. Smith and C. Mackey, Kitt Peak National Observatory; and J. Heise and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, report: “Further optical photometry of the variable reported by Bond (IAUC 6654) between May 12.00 and 12.06 UT with the 4.2-m William Herschel Telescope at La Palma shows the following magnitudes: U = 20.9 ± 0.2, B = 20.9 ± 0.3, V = 20.6 ± 0.1, R = 20.2 ± 0.1, I = 19.7 ± 0.2. Corrigenda to IAUC 6655: on May 10.0, I = 20.2 ± 0.2; on May 11.0, I = 19.1 ± 0.3 and B = 20.3 ± 0.3.”

1997 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green